Peroxyacyl nitrates contain two relatively weak bonds that can be broken by low-energy photons. The weakest is the N−O bond in the nitrate functional group of the molecule with a standard bond-dissociation energy of 201 kJ mol−1. What is the longest wavelength of a photon, in nanometres, that has enough energy to cause the photodissociation of the weakest bond. Assume all energy of the photon is efficiently used for the photodissociation.

Express your answer to three significant figures and include the appropriate units.
Answer #1

Given:

Energy of 1 mol = 2.01*10^2 KJ/mol

= 2.01*10^5 J/mol

Find energy of 1 photon first

Energy of 1 photon = energy of 1 mol/Avogadro's number

= 2.01*10^5/(6.022*10^23)

= 3.338*10^-19 J

This is energy of 1 photon

use:

E = h*c/lambda

3.338*10^-19J =(6.626*10^-34 J.s)*(3.0*10^8 m/s)/lambda

lambda = 5.955*10^-7 m

Answer: 5.96*10^-7 m